### v1

# 3Agenzia nazionale per le nuove tecnologie, l’energia e lo sviluppo economico sostenibile (ENEA), Italy 4Istituto Nazionale di Fisica Nucleare, Sezione di Bologna, Italy 5European Organization for Nuclear Research (CERN), Switzerland

Abstract
The isotopic abundances of r-process elements in the solar system are traditionally derived as residuals from the subtraction of s-process contributions from total solar abundances. However, the uncertainties in s-process nucleosynthesis—particularly those arising from Maxwellian Averaged Cross Sections (MACS)—propagate directly into the r-process residuals, affecting their reliability. Building upon the seminal work of Goriely (1999), who introduced a multi-event s-process model to quantify these uncertainties, we revisit the problem using a simplified yet effective approach. By assuming that the relative uncertainty in s-process isotopic abundances scales linearly with the MACS uncertainties from data libraries (KADoNiS), we identify a subset of isotopes for which the r-process residuals remain significantly uncertain. Using updated solar abundances (Lodders 2025) and s-process contributions from Bisterzo et al. (2014), we present a short list of isotopes that are prime candidates for improved ($n,\gamma$) measurements at CERN n_TOF in the near future. Our analysis provides a practical framework for prioritizing future experimental efforts that will profit from upgrades and enhancements of the n_TOF facility. It also highlights the need to revisit key neutron-capture cross sections to refine our understanding of the r-process isotopic abundance pattern, commonly used as a benchmark in stellar models of explosive nucleosynthesis.
pacs:
PACS-keydiscribing text of that key and PACS-keydiscribing text of that key
1 Introduction
Understanding the origin of elements heavier than iron remains one of the fundamental challenges in nuclear astrophysics. These elements are synthesized primarily through two neutron-capture processes: the slow neutron-capture process (s-process) occurring in asymptotic giant branch (AGB) stars Gallino98; Kaeppeler11; Lugaro23, and the rapid neutron-capture process (r-process) taking place in explosive environments such as neutron star mergers and core-collapse supernovae Arcones23. The traditional approach to determining the r-process contribution to solar system abundances relies on subtracting the s-process component from the total solar abundances. However, the accuracy of these r-process residuals is fundamentally limited by uncertainties in s-process nucleosynthesis models, which in turn depend critically on the Maxwellian-averaged neutron capture cross sections (MACS) of stable isotopes along the s-process path.
The seminal work of Goriely Goriely99 introduced a multi-event s-process model to quantify how uncertainties in nuclear reaction rates propagate into s-process abundance predictions and, consequently, into r-process residuals. Building on this framework and incorporating the updated s-process calculations of Bisterzo14, recent galactic chemical evolution models have provided refined estimates of s-process contributions across the nuclear chart. Nevertheless, significant uncertainties remain for several key isotopes, where large MACS uncertainties translate directly into poorly constrained r-process abundances. These uncertainties limit our ability to test theoretical models of r-process nucleosynthesis and to understand the astrophysical sites responsible for the production of the heaviest elements.
Over two decades, the CERN n_TOF facility has demonstrated its capability to measure neutron capture cross sections with high accuracy across the energy range relevant for stellar nucleosynthesis. After the CERN Long Shutdown 3 (LS3), which includes various facility upgrades, n_TOF will offer enhanced opportunities for measurements of isotopes that are currently limiting the precision of r-process abundance determinations. This paper identifies a prioritized list of isotopes for which improved $(n,\gamma)$ measurements at n_TOF would significantly reduce uncertainties in r-process residuals.
Although the primary focus of this work is on improving the accuracy of r-process isotopic abundances, we also discuss relevant aspects of s-process nucleosynthesis for each selected isotope. Many of these isotopes are located at critical positions along the s-process path—such as branching points or neutron magic numbers—where accurate cross section measurements not only reduce r-process uncertainties but also provide important constraints on s-process physics, including neutron densities, temperatures, and mixing processes in AGB stars.
From a technical perspective, the neutron-capture measurements at n_TOF employ C6D6 liquid scintillation detectors in combination with the Pulse Height Weighting Technique (PHWT) Moxon63; Macklin67. This method, which requires low-efficiency $\gamma$-ray detectors and detailed Monte Carlo simulations of the experimental setup to derive energy-dependent weighting functions, has been extensively validated at n_TOF. Dedicated systematic studies Abbondanno04; Domingo04 have demonstrated that the PHWT, when applied with detailed Monte Carlo (MC)-based weighting functions tailored to each specific sample-detector configuration, can achieve reaction yield determinations with an accuracy of 2% or better. This validation, performed using samples of various dimensions and compositions, confirmed that the technique properly accounts for the influence of sample geometry and surrounding materials on the detector response, which is a critical requirement for high-precision cross-section measurements. These developments now enable systematic cross-section determinations at n_TOF with enhanced accuracy, directly addressing the need for improved nuclear data to constrain r-process residuals and refining s-process models of nucleosynthesis Domingo23b; Domingo25.
2 Methodology
Our approach to identifying high-priority isotopes for neutron capture measurements is based on a simplified yet effective propagation of nuclear data uncertainties to r-process residuals. To this aim we utilize three primary data sources:
  • MACS values and uncertainties: The Karlsruhe Astrophysical Database of Nucleosynthesis in Stars (KADoNiS) version 1.0 kadonisv1, which provides recommended Maxwellian-averaged cross sections at $kT=30$ keV along with their uncertainties for isotopes relevant to s-process nucleosynthesis.
  • Solar system abundances: The updated solar abundance compilation by Lodders et al. (2015) Lodders2025.
  • S-process contributions: The galactic chemical evolution calculations of Bisterzo et al. (2014) Bisterzo14, which provide predictions for the s-process contribution to each isotope’s solar abundance based on detailed AGB stellar models with updated 13C-pocket structures and neutron capture networks.
The r-process residual abundance for a given isotope is obtained by subtracting the s-process contribution from the total solar abundance:
$N_{\rm r}=N_{\odot}-N_{\rm s}$ (1)
where $N_{\odot}$ is the solar system abundance and $N_{\rm s}$ is the s-process contribution predicted by stellar models.
Figure 1: Isotopic abundances in the solar system (normalized to Si $10^{6}$) showing the decomposition into s-process and r-process components. Upper panel: Total solar system abundances from Lodders2025 (open circles), s-process contributions from Bisterzo14 (blue filled circles), and derived r-process residuals $N_{\rm r}=N_{\odot}-N_{\rm s}$ (red filled circles). Lower panel: Relative s-process contributions (percentage of total solar abundance). Error bars on r-process residuals reflect the propagated uncertainties from MACS values as described in Section 2.
To estimate the uncertainty in the r-process residual, we employ a linear propagation approximation. We assume that the relative uncertainty in the s-process abundance for a given isotope scales linearly with the relative uncertainty in its MACS:
$\frac{\Delta N_{\rm s}}{N_{\rm s}}\approx\frac{\Delta\sigma}{\langle\sigma\rangle}$ (2)
where $\Delta\sigma/\langle\sigma\rangle$ is the relative MACS uncertainty from KADoNiS 1.0. Although this approximation neglects second-order effects and correlations between isotopes in network calculations, it provides a practical first-order estimate for identifying the most problematic cases. The approximation is well justified by the s-process equilibrium conditions in AGB stars: in the mass range A=90A=90–200, the product $N_{s}\langle\sigma\rangle$ remains approximately constant Kaeppeler11, which directly supports the assumed linear scaling between MACS uncertainties and s-process abundance uncertainties.
Given that uncertainties in solar system abundances from Lodders2025 are typically small ($\sim$1–3%) compared to the dominant MACS uncertainties (ranging from $\sim$10% to >>50% for many isotopes), we approximate the uncertainty in the r-process residual as:
$\Delta N_{\rm r}\approx\Delta N_{\rm s}=N_{\rm s}\cdot\frac{\Delta\sigma}{\langle\sigma\rangle}$ (3)
This simplified approach allows us to directly identify isotopes where improved MACS measurements would have the greatest impact on constraining r-process yields through reduction of uncertainties in the derived r-process residuals. The resulting r-process abundances are shown in Fig.1. As it can be inferred from Eq.3, r-process residual uncertainties become larger the higher the s-process contribution (Ns) and the higher the relative uncertainty on the MACS.
For this study we identify isotopes as high-priority targets for improved $(n,\gamma)$ measurements based on two criteria:
  1. 1.
    The relative uncertainty in the MACS from KADoNiS 1.0 must be $\geq 9\%$
  2. 2.
    The resulting relative uncertainty in the r-process residual must be $\geq 20\%$
The first criterion ensures that the isotope’s MACS is not already well-constrained by existing measurements, while the second criterion guarantees that the uncertainty propagates significantly into the r-process abundance determination. Isotopes meeting both criteria represent cases where improved cross section measurements would have the greatest impact on refining our understanding of r-process nucleosynthesis.
Table 1 lists the isotopes satisfying these criteria, showing their r-process residual abundances ($N_{\rm r}$), absolute uncertainties ($\Delta N_{\rm r}$), and the relative uncertainties in both the r-process residuals and the MACS values. In the following sections, we discuss each of these isotopes in detail, examining the current status of their cross section measurements, the astrophysical implications of the uncertainties, and the prospects for improved measurements at the CERN n_TOF facility.
[…]
97Mo $1.4(4)\times 10^{-1}$ 27 15
99Ru $1.8(6)\times 10^{-1}$ 32 25
122Sn $1.1(6)\times 10^{-1}$ 54 35
139La $1.1(5)\times 10^{-1}$ 39 10
168Er $4.9(13)\times 10^{-2}$ 27 18
184W $9.3(53)\times 10^{-3}$ 57 12
200Hg $4.3(9)\times 10^{-2}$ 22 10
Table 1: Isotope data with Nr values, uncertainties, and relative uncertainties for Nr and MACS.
3 Case Studies of Key Isotopes
97Mo
Molybdenum isotopes serve as critical tracers of s-process nucleosynthesis in presolar silicon carbide grains from AGB stars Stephan2019. While 96Mo is especially valuable as an s-process-only isotope, shielded from r-process production by 96Zr, 97Mo is a mixed s- and r-process isotope contributing about 47% to the solar inventory via s-process. The high-precision measurements of these isotopes in mainstream SiC grains Liu24 reveal that their compositions follow clear mixing lines between a pure s-process progenitor and solar material, with 96Mo and 97Mo abundances closely following the inverse of their neutron-capture cross sections. This behavior shows remarkably constant s-process ratios across different SiC grains, indicating that the branching at 95Zr remains minimal even when the 22Ne($\alpha$,n)25Mg reaction is marginally activated during thermal pulses at peak neutron densities of $\sim 10^{10}$ cm-3 Stephan2019; Lugaro03; Liu2014. This weak branching is sufficient to produce the observed 96Zr abundances without significantly bypassing 96Mo in favor of 97Mo and 98Mo. The observed variations in other Mo isotopes like 94Mo, 95Mo, and 100Mo provide unique constraints on neutron densities, temperatures, and timescales during s-process nucleosynthesis in the parent AGB stars.
Recent measurements of the 96Mo neutron capture cross section have been performed at the n_TOF facility (CERN) and complementary transmission measurements at GELINA (JRC-Geel) Mucciola22; Mucciola23. The measurements utilized isotopically enriched 96Mo samples with 95.90% enrichment. Capture cross sections were measured at both the EAR2 (18 m from spallation target) and EAR1 (184 m) experimental stations of n_TOF during the 2021–2022 campaigns, using arrays of C6D6 liquid scintillation detectors and newly developed sTED detectors Alcayne24; Balibrea25. Complementary transmission measurements were performed at the 10 m station of GELINA using a 6Li-glass neutron detector.
Preliminary results show improved accuracy and consistency compared to deviations existing nuclear data libraries (JENDL-3, ENDF/B, JEFF), particularly for resonance parameters up to 5 keV. A simultaneous resonance shape analysis combining the capture and transmission data is in progress to provide final recommended resonance parameters for 96Mo. These measurements are expected to significantly reduce the current 14% MACS uncertainty.
For 97Mo, no experimental neutron-capture cross section measurement exists to date and an experiment proposal has been approved for a high-resolution high-accuracy measurement at CERN n_TOF Busso2024. Given the similar nuclear structure and expected cross section magnitude compared to 96Mo, a measurement achieving comparable precision (reducing the current 15% MACS uncertainty kadonisv1 to $\sim$4–5%) is feasible. This would substantially improve the determination of the r-process residual, currently uncertain by $\sim$27%.
99Ru
[…]
Despite its importance as the decay product of this crucial branching point, 99Ru represents a significant gap in our nuclear data. Although the relative uncertainty quoted for the MACS of this nucleus in KADoNiS is only 15%, the recommended neutron-capture cross section is purely based on theoretical estimates. Given that theoretical predictions more typically carry uncertainties as high as a factor of 2 even for nuclei close to stability, the quoted 15% uncertainty is almost certainly underestimated. Moreover, this is one of the very few stable isotopes where no direct neutron-capture measurement exists in the stellar energy range. The apparent moderate uncertainty on the r-process residual (32%) may be a consequence of the optimistic uncertainty assigned to the unknown neutron-capture cross section. With a natural abundance of 12.76% and a large expected cross section of 631 mb at $kT=30$ keV, producing an enriched sample of 99Ru for a time-of-flight measurement should be straightforward. Reducing the uncertainty to the 4-5% level should lead to an r-process residual accurate to 6%.
From the viewpoint of the s-process, this measurement would be extremely valuable for a fully consistent interpretation of the first detection of extinct 99Tc in presolar silicon carbide grains, which was achieved by measuring an anomaly in its decay product 99Ru using resonance ionization mass spectrometry Savina2004. The isotopic composition of ruthenium in individual SiC stardust grains bears the signature of s-process nucleosynthesis in asymptotic giant branch stars, with an enrichment in 99Ru explained by the in situ decay of 99Tc that was present when the grains formed. This finding directly connects to Merrill’s 1952 observation of technetium spectral lines in S-type stars Merrill52, which provided the first evidence that elements are synthesized within stars. By detecting the radiogenic 99Ru signature left by now-extinct 99Tc, Savina et al. Savina2004 confirmed that the majority of presolar SiC grains originated from low-mass AGB stars and established that the amount of 99Tc produced in these stars is insufficient to leave a detectable 99Ru anomaly in bulk early solar system materials. This work allowed stellar nucleosynthesis models to be quantitatively tested with unprecedented accuracy by bridging spectroscopic observations of active stellar nucleosynthesis with laboratory measurements of preserved stellar ejecta.
122Sn
The pioneering work of Macklin and co-workers in 1962 Macklin62 demonstrated that tin’s ten stable isotopes provide exceptional opportunities to disentangle the contributions of different neutron-capture processes. According to the original hypothesis of heavy-element nucleosynthesis BBFH, certain tin isotopes can only be produced through specific pathways: 116Sn cannot form via r-process, while 122Sn and 124Sn were thought to be exclusive products of the rapid neutron capture process. By measuring neutron capture cross-sections of the tin isotopes at stellar temperatures and examining the inverse relationship between these cross-sections and isotopic abundances it was possible to quantify for the first time the relative contributions of s-process versus r-process nucleosynthesis in the solar system’s material Macklin62. This approach confirmed theoretical predictions about how these processes operate in different stellar environments: the s-process occurring in giant stars and the r-process in explosive environments. The approximate constancy of the product of neutron capture cross-section and abundance for s-process isotopes provided direct experimental validation of nucleosynthesis models, establishing tin as an ideal “laboratory” element for testing our understanding of how heavy elements form in stars.
Today, thanks to detailed AGB models we know that 122Sn is not an exclusive product of the r-process Bisterzo14. However, despite the critical role of 122Sn in disentangling s- and r-process contributions in this mass region, experimental constraints on its neutron capture cross section remain limited. The current nuclear data situation relies primarily on a single time-of-flight measurement by Timokhov1989 and two activation measurements Stadler1998; Murty1973, one of them unpublished. The recommended MACS carries an uncertainty of 35%, which propagates to an r-process residual uncertainty exceeding 50% when performing classical s/r-process decomposition. This large uncertainty significantly hampers our ability to use 122Sn as a diagnostic tool for understanding the branching-point physics and for constraining galactic chemical evolution models Bisterzo14.
Given its relatively high natural abundance of 4.63%, 122Sn represents an excellent candidate for high-precision measurements combining time-of-flight and activation techniques. New measurements at the CERN n_TOF facility, utilizing both the EAR1 and EAR2 beam lines as well as the NEAR station for activation studies Patronis25, would provide the complementary data needed to substantially reduce these uncertainties and better constrain the s-process contribution to this key branching-point isotope.
139La
139La is the most abundant isotope of lanthanum (>99.9>99.9%) and is primarily produced in the s-process, with an estimated contribution of 75% Bisterzo14. It is of particular interest since it has a magic neutron number ($N=82$) and as a consequence a small neutron capture cross section compared to neighbouring isotopes. Hence, the 139La($n,\gamma$) reaction represents a bottleneck in the s-process flow where matter accumulates. Based on the current uncertainties of the Maxwellian averaged cross section of around 10%, the resultant uncertainty of the solar r-process abundance is around 39%.
There are several measurements of the 139La($n,\gamma$) cross section kadonisv1. While available experimental MACS values at $kT=30$ keV range from 22.2 to 50 mb Dillmann08, the three most recent results quote consistent values of $31.6\pm 0.8$ mb Obrien03, $31.2$ mb (renormalised from a measurement at $kT=5$ keV) Winckler_2006 and $32.4\pm 2.1$ mb Terlizzi07. The recommended value in the KADoNiS database and hence the value most frequently used in stellar model calculations is the result from measurement at n_TOF by Terlizzi et al Terlizzi07. This measurement covered neutron energy ranges from 0.6 eV to 9 keV. The Maxwellian averaged cross sections were calculated by using the extracted resonance information combined with the evaluated cross sections from the JENDL-3.3 database between 9 keV and 1 MeV, which has been scaled to match the experimental cross section at lower neutron energy. The total uncertainty of the MACS at 30 keV is dominated by the uncertainty assumed on the JENDL-3.3 cross section (10%). The activation measurement by O’Brien et al. Obrien03 is the only direct measurement at 30 keV neutron energy. The quoted uncertainty is small and does not include the uncertainty of the 197Au($n,\gamma$) the spectrum averaged cross section used as a reference reaction. Incidentally, the 197Au($n,\gamma$) in the keV neutron energy region has recently been updated Lederer11; Massimi14; Reifarth2018, which would lead to a small reduction of the value obtained by Obrien03.
Concerning the n_TOF measurement Terlizzi07, the energy range analysed was limited by the increase of in-beam $\gamma$-ray background in the keV neutron energy range, combined with low statistics due to the small neutron capture cross section. Since then, the spallation target at n_TOF was replaced, and now borated water is used as neutron moderator. This results in a significant reduction of the in-beam $\gamma$-ray background (see Fig.13 in Ref.Massimi10, and Refs. Lerendegui16a; Lerendegui16b), and neutron capture cross sections have since been routinely measured up to tens or hundreds of keV, even for comparably small cross sections, e.g. Lederer14; Gawlik21; Amaducci24. A new measurement, aiming for high counting statistics at n_TOF would allow to access the entire neutron energy range relevant for AGB stars (the main producers of the s-process component of 139La), and reduce uncertainties to 5-6%, typically achieved at n_TOF. This can be combined with an activation measurement at n_TOF NEAR. In parallel, a re-normalisation of the O’Brien Obrien03 values based on the updated 197Au($n,\gamma$) cross section will help to provide a second independent value for the MACS at $kT=30$ keV.
168Er
Erbium isotopes, including 168Er, occupy a critical region in the landscape of s-process nucleosynthesis. Located beyond the N = 82 neutron magic number in the mass range A $\sim$ 166–170, the erbium isotopes lie in the transition zone between the second s-process abundance peak (A $\sim$ 140) and the third peak (A $\sim$ 208). The s-process dominates the production of elements in the erbium-ytterbium (Er–Yb) region, with contributions of 70–90% of the solar system abundances. This mass region is characterized by deformed, mid-shell nuclei with low-lying excited states that can significantly impact stellar neutron capture rates through thermal population effects. The relatively large uncertainties in the MACS for 168Er represents a limiting factor in the precision ($\sim$18%) of s-process low-mass AGB models, which propagates to almost 30% for the corresponding r-process component. With 168Er having a relatively large cross section (¿300 mb at 30keV) and a natural abundance of approximately 27%, a new TOF measurement at n_TOF EAR1 or EAR2 should be able to deliver new MACS values with an uncertainty down to $\sim$5% in the 30 keV energy range.
184W
[…]
The most recent direct measurement Marganiec09 employed the activation technique at the Karlsruhe Van de Graaff accelerator, yielding a MACS at $kT=30$ keV of $221\pm 12$ mb (5.4% uncertainty). This value confirmed the previously recommended MACS of $224\pm 10$ mb Bao00 and showed good agreement with earlier time-of-flight measurements Macklin83; Beer82. However, when combined with new data for the neighboring isotopes 180W and 186W, the measurements revealed a significantly flatter cross-section trend with mass number than predicted by Hauser-Feshbach calculations Rauscher00; Goriely, highlighting remaining challenges in theoretical modeling and indicating that statistical model predictions commonly used for r-process network calculations may overestimate cross sections for the neutron-rich tungsten isotopes.
From the viewpoint of enhanced accuracy measurements, improving the presently recommended relative uncertainty of $\sim$12%kadonisv1 down to 4% could still help to reduce the 57% uncertainty on the r-process isotopic abundance down to $\sim$20%. Such a measurement at CERN n_TOF could benefit of the availability of both TOF- and activation techniques and new instrumentation at EAR1, EAR2 Alcayne24; Balibrea25 and NEAR Patronis25 stations. A measurement which, considering the relatively large cross section and natural abundance of 184W (30%) should be rather straight-forward.
200Hg
The r-process residual derived from 200Hg may serve as an important anchor point for constraining the r-process abundance distribution in the $A\sim 200$ mass region. Together with the other mercury isotopes (199,201,202Hg), the 200Hg data help to define the relatively flat r-process abundance plateau expected in this mass range, lying between the $N=126$ r-process peak and the lead region. This information is helpful for validating r-process nucleosynthesis models and for understanding the conditions under which rapid neutron capture occurred in the early Galaxy. With a measured MACS of $115\pm 12$ mb at $kT=30$ keV, 200Hg represents a key (s+r) process isotope that receives rather balanced contributions from both slow and rapid neutron-capture processes. Thus, the accurate cross section determination enables quantitative decomposition of the 200Hg solar abundance into its s- and r-process components. Using the measured cross section and the $\sigma N$ systematics of the s-process, Beer and Macklin determined that of the total solar abundance of $(0.045\pm 0.007)/10^{6}$ Si, the s-process contributes 42(5)% while the r-process accounts for 55(16)%Beer85. A new measurements could significantly decrease the uncertainty, down to $\sim$5%, thus lowering also the relative r-process uncertainty down to the level of $\sim$10%.
4 Summary and conclusions
We have presented a systematic analysis of neutron capture cross section uncertainties and their impact on r-process residual determinations for isotopes in the mass range $A\sim 96$–200. By combining MACS data from KADoNiS 1.0 kadonisv1, solar system abundances from Lodders2025, and s-process predictions from Bisterzo14, we have identified a set of high-priority targets where improved measurements would significantly reduce uncertainties in derived r-process yields.
Our analysis reveals that seven isotopes meet the dual criteria of MACS relative uncertainty $\geq 9\%$ and r-process residual relative uncertainty $\geq 20\%$: 97Mo, 99Ru, 122Sn, 139La, 168Er, 184W and 200Hg. These isotopes span critical regions of the chart of nuclides, including the $N=50$ shell closure (97Mo), s-process branching points at A=122A=122–123 (122Sn), the neutron magic number $N=82$ (139La), and the rare earth region where main s-process contributions dominate (168Er). The case of 99Ru is particularly noteworthy, as current KADoNiS recommendations rely entirely on theoretical calculations with no direct experimental constraints.
The enhanced capabilities of CERN n_TOF following Long Shutdown 3 present an excellent opportunity to address these measurement needs. The combination of the high-flux EAR1 and EAR2 Pavon25 beam lines with the newly commissioned NEAR station Patronis25 provides optimal conditions for both time-of-flight and activation measurements across the relevant energy range. The improved neutron flux, enhanced detection systems Alcayne24; Balibrea25, and extended measurement capabilities will enable precision determinations of MACS values with target uncertainties below 5% for most of the identified isotopes.
Beyond the immediate goal of constraining r-process residuals, these measurements will provide critical input for s-process nucleosynthesis models themselves. Isotopes such as 122Sn and 139La are located at key branching points and bottlenecks in the s-process path, where neutron capture rates directly influence the final abundance distributions produced in AGB stars. Improved cross sections for these nuclei will therefore refine our understanding of both the weak and main s-process components, with implications for galactic chemical evolution models and the interpretation of abundance patterns in metal-poor stars.
The linear uncertainty propagation approach employed in this work, while simplified, provides a practical framework for prioritizing experimental efforts. Future refinements could incorporate full network calculations to account for correlation effects and non-linear sensitivities, particularly for branching point isotopes where competition between neutron capture and $\beta$-decay determines the s-process flow. Nevertheless, our first-order analysis successfully identifies the most critical cases where nuclear data improvements will have maximum impact.
In conclusion, the post-LS3 experimental program at CERN n_TOF represents a significant step forward in addressing long-standing uncertainties in neutron capture cross sections that limit our ability to cleanly separate s-process and r-process contributions to solar system abundances. The measurements proposed here will not only improve constraints on r-process nucleosynthesis yields but will also advance our understanding of s-process physics in AGB stars, contributing to a more complete picture of heavy element production in the Galaxy.
Acknowledgments
We dedicate this work to the memory of Roberto Gallino, whose pioneering theoretical work on s-process nucleosynthesis in AGB stars provided the essential framework for interpreting neutron-capture measurements such as those performed at FZK Karlsruhe and CERN n_TOF Gallino23. His detailed stellar models, incorporating the physics of 13C-pockets and nuclear-physics networks, have been instrumental in calculating the s-process contributions needed to understand the composition of our solar system and the chemical evolution of our galaxy. His legacy continues to guide our efforts to constrain the origins of the heavy elements in the cosmos.
CDP acknowledges fruitful discussions with Almudena Arcones and Gabriel Martínez Pinedo at the EMMI Workshop ”Nucleosynthesis of heavy elements: r-process” held in Hirschegg, Austria, in 2025. The authors acknowledge K. Lodders and I. Dillmann for facilitating quick access to the data of their publications and evaluations.
The authors acknowledge support from all the funding agencies of participating institutions. Part of this work was supported by the MCIN/AEI 10.13039/501100011033 under grants Severo Ochoa CEX2023-001292-S and PID2022-138297NB-C21 and UKRI STFC Grant ST/Y000293/1. The corresponding author has applied a Creative Commons Attribution (CC BY-NC-ND 4.0) license.
Supplemental material
[…]
Isotope $N_{\odot}$ $N_{r}$ $N_{r}/N_{\odot}$ $N_{s}$ $N_{s}/N_{\odot}$ $\Delta\sigma/\langle\sigma\rangle$
$(\times 10^{6}~\mathrm{Si})$ $(\times 10^{6}~\mathrm{Si})$ (%) $(\times 10^{6}~\mathrm{Si})$ (%) (%)
80Kr $1.34$ $1.16(7)$ 87 $1.78(73)\times 10^{-1}$ 13 5
82Kr $6.70$ $4.86(45)$ 73 $1.84(45)$ 27 7
83Kr $6.63$ $6.00(35)$ 90 $6.30(346)\times 10^{-1}$ 10 5
84Kr $3.26\times 10^{1}$ $2.88(25)\times 10^{1}$ 88 $3.81(246)$ 12 8
86Kr $9.82$ $8.33(58)$ 85 $1.49(58)$ 15 6
85Rb $5.02$ $4.25(19)$ 85 $7.68(188)\times 10^{-1}$ 15 4
87Rb $2.07$ $1.55(11)$ 75 $5.15(110)\times 10^{-1}$ 25 5
86Sr $2.37$ $9.22(178)\times 10^{-1}$ 39 $1.45(18)$ 61 8
87Sr $1.66$ $6.94(67)\times 10^{-1}$ 42 $9.66(67)\times 10^{-1}$ 58 4
88Sr $1.98\times 10^{1}$ $5.72(63)$ 29 $1.41(6)\times 10^{1}$ 71 3
89Y $4.49$ $1.26(13)$ 28 $3.23(13)$ 72 3
90Zr $5.41$ $2.15(25)$ 40 $3.26(25)$ 60 5
91Zr $1.18$ $3.40(75)\times 10^{-1}$ 29 $8.40(75)\times 10^{-1}$ 71 6
92Zr $1.80$ $5.72(143)\times 10^{-1}$ 32 $1.23(14)$ 68 8
94Zr $1.82$ $2.98(69)\times 10^{-1}$ 16 $1.52(7)$ 84 4
96Zr $2.90\times 10^{-1}$ $1.78(18)\times 10^{-1}$ 61 $1.12(18)\times 10^{-1}$ 39 6
93Nb $7.66\times 10^{-1}$ $3.37(15)\times 10^{-1}$ 44 $4.29(15)\times 10^{-1}$ 56 2
94Mo $2.38\times 10^{-1}$ $2.36(32)\times 10^{-1}$ 99 $2.38(3216)\times 10^{-3}$ 1 14
95Mo $4.12\times 10^{-1}$ $2.25(43)\times 10^{-1}$ 55 $1.87(43)\times 10^{-1}$ 45 10
96Mo $4.33\times 10^{-1}$ $9.44(619)\times 10^{-2}$ 22 $3.39(62)\times 10^{-1}$ 78 14
97Mo $2.49\times 10^{-1}$ $1.41(38)\times 10^{-1}$ 57 $1.08(38)\times 10^{-1}$ 43 15
98Mo $6.30\times 10^{-1}$ $2.68(42)\times 10^{-1}$ 43 $3.62(42)\times 10^{-1}$ 58 7
100Mo $2.53\times 10^{-1}$ $2.47(18)\times 10^{-1}$ 98 $5.82(1788)\times 10^{-3}$ 2 7
99Ru $2.26\times 10^{-1}$ $1.79(57)\times 10^{-1}$ 79 $4.75(566)\times 10^{-2}$ 21 25
100Ru $2.23\times 10^{-1}$ $4.44(147)\times 10^{-2}$ 20 $1.79(15)\times 10^{-1}$ 80 7
101Ru $3.02\times 10^{-1}$ $2.63(20)\times 10^{-1}$ 87 $3.87(199)\times 10^{-2}$ 13 7
102Ru $5.58\times 10^{-1}$ $3.16(36)\times 10^{-1}$ 57 $2.42(36)\times 10^{-1}$ 43 6
104Ru $3.29\times 10^{-1}$ $3.24(16)\times 10^{-1}$ 99 $4.61(1625)\times 10^{-3}$ 1 5
103Rh $3.41\times 10^{-1}$ $3.01(6)\times 10^{-1}$ 88 $4.02(57)\times 10^{-2}$ 12 2
104Pd $1.54\times 10^{-1}$ $2.48(33)\times 10^{-2}$ 16 $1.29(3)\times 10^{-1}$ 84 2
105Pd $3.09\times 10^{-1}$ $2.76(5)\times 10^{-1}$ 89 $3.31(54)\times 10^{-2}$ 11 2
106Pd $3.79\times 10^{-1}$ $2.28(38)\times 10^{-1}$ 60 $1.51(38)\times 10^{-1}$ 40 10
108Pd $3.67\times 10^{-1}$ $1.82(28)\times 10^{-1}$ 50 $1.85(28)\times 10^{-1}$ 50 8
110Pd $1.62\times 10^{-1}$ $1.59(26)\times 10^{-1}$ 98 $2.59(2570)\times 10^{-3}$ 2 16
107Ag $2.61\times 10^{-1}$ $2.61(8)\times 10^{-1}$ 100 $2.61(7879)\times 10^{-4}$ 0 3
109Ag $2.42\times 10^{-1}$ $1.87(7)\times 10^{-1}$ 77 $5.47(71)\times 10^{-2}$ 23 3
108Cd $1.40\times 10^{-2}$ $1.38(6)\times 10^{-2}$ 99 $1.54(642)\times 10^{-4}$ 1 5
110Cd $1.99\times 10^{-1}$ $4.46(35)\times 10^{-2}$ 22 $1.54(3)\times 10^{-1}$ 78 2
 [^1] Cd $2.04\times 10^{-1}$ $1.53(3)\times 10^{-1}$ 75 $5.08(34)\times 10^{-2}$ 25 2
112Cd $3.85\times 10^{-1}$ $1.93(3)\times 10^{-1}$ 50 $1.92(3)\times 10^{-1}$ 50 1
113Cd $1.95\times 10^{-1}$ $1.39(10)\times 10^{-1}$ 72 $5.56(98)\times 10^{-2}$ 28 5
114Cd $4.58\times 10^{-1}$ $1.87(5)\times 10^{-1}$ 41 $2.71(5)\times 10^{-1}$ 59 1
116Cd $1.20\times 10^{-1}$ $1.09(1)\times 10^{-1}$ 91 $1.06(14)\times 10^{-2}$ 9 1
113In $8.00\times 10^{-3}$ $8.00(112)\times 10^{-3}$ 100 0 0 14
115In $1.71\times 10^{-1}$ $1.20(18)\times 10^{-1}$ 70 $5.10(181)\times 10^{-2}$ 30 11
114Sn $2.40\times 10^{-2}$ $2.40(3)\times 10^{-2}$ 100 $2.40(3214)\times 10^{-5}$ 0 1
115Sn $1.30\times 10^{-2}$ $1.25(3)\times 10^{-2}$ 96 $5.46(331)\times 10^{-4}$ 4 3
116Sn $5.37\times 10^{-1}$ $1.70(12)\times 10^{-1}$ 32 $3.67(12)\times 10^{-1}$ 68 2
117Sn $2.83\times 10^{-1}$ $1.74(16)\times 10^{-1}$ 62 $1.09(16)\times 10^{-1}$ 38 6
118Sn $8.94\times 10^{-1}$ $3.99(9)\times 10^{-1}$ 45 $4.95(9)\times 10^{-1}$ 55 1
119Sn $3.17\times 10^{-1}$ $1.68(13)\times 10^{-1}$ 53 $1.49(13)\times 10^{-1}$ 47 4
120Sn $1.20$ $4.49(30)\times 10^{-1}$ 37 $7.54(30)\times 10^{-1}$ 63 2
122Sn $1.71\times 10^{-1}$ $1.09(59)\times 10^{-1}$ 64 $6.22(590)\times 10^{-2}$ 36 35
121Sb $1.94\times 10^{-1}$ $1.34(6)\times 10^{-1}$ 69 $5.99(58)\times 10^{-2}$ 31 3
123Sb $1.45\times 10^{-1}$ $1.38(4)\times 10^{-1}$ 95 $7.25(431)\times 10^{-3}$ 5 3
Table 3: [Continuation]
Isotope $N_{\odot}$ $N_{r}$ $N_{r}/N_{\odot}$ $N_{s}$ $N_{s}/N_{\odot}$ $\Delta\sigma/\langle\sigma\rangle$
$(\times 10^{6}~\mathrm{Si})$ $(\times 10^{6}~\mathrm{Si})$ (%) $(\times 10^{6}~\mathrm{Si})$ (%) (%)
122Te $1.24\times 10^{-1}$ $3.27(11)\times 10^{-2}$ 26 $9.13(11)\times 10^{-2}$ 74 1
123Te $4.30\times 10^{-2}$ $1.11(4)\times 10^{-2}$ 26 $3.19(4)\times 10^{-2}$ 74 1
124Te $2.29\times 10^{-1}$ $5.02(30)\times 10^{-2}$ 22 $1.79(3)\times 10^{-1}$ 78 1
125Te $3.39\times 10^{-1}$ $2.80(3)\times 10^{-1}$ 83 $5.90(31)\times 10^{-2}$ 17 1
126Te $8.99\times 10^{-1}$ $5.73(15)\times 10^{-1}$ 64 $3.26(15)\times 10^{-1}$ 36 2
128Te $1.50$ $1.45(4)$ 97 $4.96(440)\times 10^{-2}$ 3 3
127I $1.59$ $1.53(21)$ 96 $6.04(2138)\times 10^{-2}$ 4 13
128Xe $1.24\times 10^{-1}$ $2.94(17)\times 10^{-2}$ 24 $9.46(17)\times 10^{-2}$ 76 1
129Xe $1.54$ $1.49(3)$ 97 $4.30(299)\times 10^{-2}$ 3 2
130Xe $2.44\times 10^{-1}$ $3.64(39)\times 10^{-2}$ 15 $2.08(4)\times 10^{-1}$ 85 2
132Xe $1.48$ $1.08(12)$ 73 $3.96(123)\times 10^{-1}$ 27 8
134Xe $5.45\times 10^{-1}$ $5.23(49)\times 10^{-1}$ 96 $2.23(486)\times 10^{-2}$ 4 9
133Cs $3.67\times 10^{-1}$ $3.17(20)\times 10^{-1}$ 86 $4.95(205)\times 10^{-2}$ 14 6
134Ba $12\times 10^{-1} 1.1$$11(4)\times 10^{-1}$ 99 $1.01(356)\times 10^{-3}$ 1 3
135Ba $3.04\times 10^{-1}$ $2.17(10)\times 10^{-1}$ 72 $8.66(100)\times 10^{-2}$ 28 3
136Ba $3.62\times 10^{-1}$ $3.57(12)\times 10^{-1}$ 99 $4.71(1185)\times 10^{-3}$ 1 3
137Ba $5.18\times 10^{-1}$ $1.91(17)\times 10^{-1}$ 37 $3.27(17)\times 10^{-1}$ 63 3
138Ba $3.31$ $2.71(160)\times 10^{-1}$ 8 $3.04(16)$ 92 5
139La $4.68\times 10^{-1}$ $1.15(45)\times 10^{-1}$ 24 $3.53(45)\times 10^{-1}$ 76 10
140Ce $1.05$ $8.40(394)\times 10^{-2}$ 8 $9.66(39)\times 10^{-1}$ 92 4
142Ce $1.32\times 10^{-1}$ $1.06(4)\times 10^{-1}$ 80 $2.57(44)\times 10^{-2}$ 20 3
141Pr $1.79\times 10^{-1}$ $8.97(22)\times 10^{-2}$ 50 $8.93(22)\times 10^{-2}$ 50 1
142Nd $2.37\times 10^{-1}$ $4.03(473)\times 10^{-3}$ 2 $2.33(5)\times 10^{-1}$ 98 2
143Nd $1.05\times 10^{-1}$ $7.05(13)\times 10^{-2}$ 67 $45(13)\times 10^{-2} 33$ 1
144Nd $2.08\times 10^{-1}$ $9.94(38)\times 10^{-2}$ 48 $1.09(4)\times 10^{-1}$ 52 2
145Nd $7.70\times 10^{-2}$ $5.68(8)\times 10^{-2}$ 74 $2.02(8)\times 10^{-2}$ 26 1
146Nd $1.50\times 10^{-1}$ $5.10(16)\times 10^{-2}$ 34 $9.90(16)\times 10^{-2}$ 66 1
148Nd $5.00\times 10^{-2}$ $4.13(6)\times 10^{-2}$ 83 $8.65(65)\times 10^{-3}$ 17 1
148Sm $4.22\times 10^{-2}$ $4.08(11)\times 10^{-2}$ 97 $1.43(114)\times 10^{-3}$ 3 3
149Sm $3.07\times 10^{-2}$ $2.67(7)\times 10^{-2}$ 87 $3.96(74)\times 10^{-3}$ 13 2
150Sm $3.77\times 10^{-2}$ 0 0 $3.77(4)\times 10^{-2}$ 100 1
152Sm $2.01\times 10^{-2}$ $1.55(2)\times 10^{-2}$ 77 $4.58(19)\times 10^{-3}$ 23 1
154Sm $7.29\times 10^{-2}$ $7.11(33)\times 10^{-2}$ 98 $1.82(326)\times 10^{-3}$ 2 4
151Eu $4.98\times 10^{-2}$ $4.69(15)\times 10^{-2}$ 94 $2.94(148)\times 10^{-3}$ 6 3
153Eu $5.43\times 10^{-2}$ $5.10(37)\times 10^{-2}$ 94 $3.31(374)\times 10^{-3}$ 6 7
152Gd $7.10\times 10^{-4}$ $9.51(115)\times 10^{-5}$ 13 $6.15(12)\times 10^{-4}$ 87 2
154Gd $7.68\times 10^{-3}$ $8.37(90)\times 10^{-4}$ 11 $6.84(9)\times 10^{-3}$ 89 1
155Gd $5.21\times 10^{-2}$ $4.90(6)\times 10^{-2}$ 94 $3.13(59)\times 10^{-3}$ 6 1
156Gd $7.21\times 10^{-2}$ $5.92(6)\times 10^{-2}$ 82 $1.29(6)\times 10^{-2}$ 18 1
157Gd $5.51\times 10^{-2}$ $4.90(6)\times 10^{-2}$ 89 $6.12(60)\times 10^{-3}$ 11 1
158Gd $8.75\times 10^{-2}$ $6.34(8)\times 10^{-2}$ 72 $2.41(8)\times 10^{-2}$ 28 1
160Gd $7.70\times 10^{-2}$ $7.65(108)\times 10^{-2}$ 99 $5.39(10816)\times 10^{-4}$ 1 14
159Tb $6.37\times 10^{-2}$ $5.86(90)\times 10^{-2}$ 92 $5.10(904)\times 10^{-3}$ 8 14
160Dy $9.80\times 10^{-3}$ $9.80(132)\times 10^{-4}$ 10 $8.82(13)\times 10^{-3}$ 90 1
161Dy $7.93\times 10^{-2}$ $7.51(8)\times 10^{-2}$ 95 $4.20(77)\times 10^{-3}$ 5 1
162Dy $1.07\times 10^{-1}$ $8.98(10)\times 10^{-2}$ 84 $1.71(10)\times 10^{-2}$ 16 1
163Dy $1.05\times 10^{-1}$ $9.99(10)\times 10^{-2}$ 96 $4.60(103)\times 10^{-3}$ 4 1
164Dy $1.19\times 10^{-1}$ $9.03(17)\times 10^{-2}$ 76 $2.83(17)\times 10^{-2}$ 24 1
165Ho $9.06\times 10^{-2}$ $8.31(134)\times 10^{-2}$ 92 $7.52(1340)\times 10^{-3}$ 8 15
164Er $4.00\times 10^{-3}$ $6.64(169)\times 10^{-4}$ 17 $3.34(17)\times 10^{-3}$ 83 4
166Er $8.80\times 10^{-2}$ $7.33(94)\times 10^{-2}$ 83 $1.47(94)\times 10^{-2}$ 17 11
167Er $6.00\times 10^{-2}$ $5.44(96)\times 10^{-2}$ 91 $5.64(960)\times 10^{-3}$ 9 16
168Er $7.10\times 10^{-2}$ $4.87(129)\times 10^{-2}$ 69 $2.23(129)\times 10^{-2}$ 31 18
170Er $3.90\times 10^{-2}$ $3.41(16)\times 10^{-2}$ 87 $4.91(156)\times 10^{-3}$ 13 4
169Tm $4.10\times 10^{-2}$ $3.73(25)\times 10^{-2}$ 91 $3.73(250)\times 10^{-3}$ 9 6
170Yb $8.00\times 10^{-3}$ $3.36(75)\times 10^{-4}$ 4 $7.66(7)\times 10^{-3}$ 96 1
171Yb $3.60\times 10^{-2}$ $2.78(4)\times 10^{-2}$ 77 $8.17(36)\times 10^{-3}$ 23 1
172Yb $5.60\times 10^{-2}$ $3.14(5)\times 10^{-2}$ 56 $2.46(5)\times 10^{-2}$ 44 1
173Yb $4.20\times 10^{-2}$ $3.03(4)\times 10^{-2}$ 72 $1.17(4)\times 10^{-2}$ 28 1
174Yb $8.30\times 10^{-2}$ $3.25(10)\times 10^{-2}$ 39 $5.05(10)\times 10^{-2}$ 61 1
176Yb $3.40\times 10^{-2}$ $3.16(6)\times 10^{-2}$ 93 $2.45(59)\times 10^{-3}$ 7 2
Table 4: [Continuation]
Isotope $N_{\odot}$ $N_{r}$ $N_{r}/N_{\odot}$ $N_{s}$ $N_{s}/N_{\odot}$ $\Delta\sigma/\langle\sigma\rangle$
$(\times 10^{6}~\mathrm{Si})$ $(\times 10^{6}~\mathrm{Si})$ (%) $(\times 10^{6}~\mathrm{Si})$ (%) (%)
175Lu $3.73\times 10^{-2}$ $3.05(3)\times 10^{-2}$ 82 $6.75(31)\times 10^{-3}$ 18 1
176Lu $1.10\times 10^{-3}$ $1.04(1)\times 10^{-3}$ 94 $6.05(94)\times 10^{-5}$ 6 1
176Hf $8.00\times 10^{-3}$ $7.99(22)\times 10^{-3}$ 100 $8.00(21549)\times 10^{-6}$ 0 3
177Hf $2.90\times 10^{-2}$ $2.40(2)\times 10^{-2}$ 83 $5.02(23)\times 10^{-3}$ 17 1
178Hf $4.30\times 10^{-2}$ $1.83(4)\times 10^{-2}$ 42 $2.47(4)\times 10^{-2}$ 58 1
179Hf $2.10\times 10^{-2}$ $1.23(2)\times 10^{-2}$ 59 $8.65(18)\times 10^{-3}$ 41 1
180Hf $5.50\times 10^{-2}$ $5.83(67)\times 10^{-3}$ 11 $4.92(7)\times 10^{-2}$ 89 1
181Ta $2.17\times 10^{-2}$ $1.16(4)\times 10^{-2}$ 53 $1.01(4)\times 10^{-2}$ 47 2
180W $2.00\times 10^{-4}$ $1.91(20)\times 10^{-4}$ 96 $9.00(1951)\times 10^{-6}$ 4 10
182W $3.80\times 10^{-2}$ $1.38(15)\times 10^{-2}$ 36 $2.42(15)\times 10^{-2}$ 64 4
183W $2.10\times 10^{-2}$ $7.94(91)\times 10^{-3}$ 38 $1.31(9)\times 10^{-2}$ 62 4
184W $4.40\times 10^{-2}$ $9.33(529)\times 10^{-3}$ 21 $3.47(53)\times 10^{-2}$ 79 12
186W $4.10\times 10^{-2}$ $2.36(22)\times 10^{-2}$ 58 $1.74(22)\times 10^{-2}$ 42 5
185Re $2.11\times 10^{-2}$ $1.54(9)\times 10^{-2}$ 73 $5.70(85)\times 10^{-3}$ 27 4
187Re $3.80\times 10^{-2}$ $3.44(19)\times 10^{-2}$ 91 $3.57(193)\times 10^{-3}$ 9 5
186Os $1.10\times 10^{-2}$ $1.06(5)\times 10^{-2}$ 96 $3.85(452)\times 10^{-4}$ 4 4
187Os $9.00\times 10^{-3}$ $5.63(30)\times 10^{-3}$ 63 $3.37(30)\times 10^{-3}$ 37 3
188Os $9.00\times 10^{-2}$ $6.46(43)\times 10^{-2}$ 72 $2.54(43)\times 10^{-2}$ 28 5
189Os $1.10\times 10^{-1}$ $1.05(3)\times 10^{-1}$ 95 $5.28(305)\times 10^{-3}$ 5 3
190Os $1.79\times 10^{-1}$ $1.53(7)\times 10^{-1}$ 85 $2.61(71)\times 10^{-2}$ 15 4
192Os $2.78\times 10^{-1}$ $2.69(12)\times 10^{-1}$ 97 $9.17(1216)\times 10^{-3}$ 3 4
191Ir $32\times 10^{-1} 2.2$$28(7)\times 10^{-1}$ 98 $4.41(739)\times 10^{-3}$ 2 3
193Ir $3.91\times 10^{-1}$ $3.86(28)\times 10^{-1}$ 99 $5.47(2754)\times 10^{-3}$ 1 7
192Pt $10.00\times 10^{-3}$ $1.88(41)\times 10^{-3}$ 19 $8.12(41)\times 10^{-3}$ 81 4
194Pt $4.00\times 10^{-1}$ $3.80(11)\times 10^{-1}$ 95 $1.96(113)\times 10^{-2}$ 5 3
195Pt $4.12\times 10^{-1}$ $4.04(12)\times 10^{-1}$ 98 $8.24(1216)\times 10^{-3}$ 2 3
196Pt $3.08\times 10^{-1}$ $2.70(10)\times 10^{-1}$ 88 $3.79(96)\times 10^{-2}$ 12 3
198Pt $90\times 10^{-2} 8.8$$89(38)\times 10^{-2}$ 100 $8.90(37872)\times 10^{-5}$ 0 4
197Au $2.01\times 10^{-1}$ $1.89(2)\times 10^{-1}$ 94 $1.23(23)\times 10^{-2}$ 6 1
198Hg $3.90\times 10^{-2}$ $6.71(338)\times 10^{-3}$ 17 $3.23(34)\times 10^{-2}$ 83 9
199Hg $6.50\times 10^{-2}$ $5.10(40)\times 10^{-2}$ 78 $1.40(40)\times 10^{-2}$ 22 6
200Hg $8.90\times 10^{-2}$ $4.28(93)\times 10^{-2}$ 48 $4.62(93)\times 10^{-2}$ 52 10
201Hg $5.10\times 10^{-2}$ $3.07(27)\times 10^{-2}$ 60 $2.03(27)\times 10^{-2}$ 40 5
202Hg $1.14\times 10^{-1}$ $3.37(61)\times 10^{-2}$ 30 $8.03(61)\times 10^{-2}$ 70 5
204Hg $2.60\times 10^{-2}$ $2.39(25)\times 10^{-2}$ 92 $2.13(248)\times 10^{-3}$ 8 10
203Tl $5.40\times 10^{-2}$ $1.04(35)\times 10^{-2}$ 19 $4.36(35)\times 10^{-2}$ 81 7
205Tl $1.28\times 10^{-1}$ $4.26(95)\times 10^{-2}$ 33 $8.54(95)\times 10^{-2}$ 67 7
204Pb $6.60\times 10^{-2}$ $8.58(252)\times 10^{-3}$ 13 $5.74(25)\times 10^{-2}$ 87 4
206Pb $6.12\times 10^{-1}$ $1.66(12)\times 10^{-1}$ 27 $4.46(12)\times 10^{-1}$ 73 2
207Pb $6.77\times 10^{-1}$ $2.02(39)\times 10^{-1}$ 30 $4.75(39)\times 10^{-1}$ 70 6
208Pb $1.94$ $4.46(2062)\times 10^{-2}$ 2 $1.89(21)$ 98 11
209Bi $42\times 10^{-1} 1.1$$13(22)\times 10^{-1}$ 80 $2.90(223)\times 10^{-2}$ 20 16

### v2

@@
 1 Introduction
-Understanding the origin of elements heavier than iron remains one of the fundamental challenges in nuclear astrophysics.
-These elements are synthesized primarily through two neutron-capture processes: the slow neutron-capture process (s-process) occurring in asymptotic giant branch (AGB) stars Gallino98; Kaeppeler11; Lugaro23, and the rapid neutron-capture process (r-process) taking place in explosive environments such as neutron star mergers and core-collapse supernovae Arcones23.
-The traditional approach to determining the r-process contribution to solar system abundances relies on subtracting the s-process component from the total solar abundances.
-However, the accuracy of these r-process residuals is fundamentally limited by uncertainties in s-process nucleosynthesis models, which in turn depend critically on the Maxwellian-averaged neutron capture cross sections (MACS) of stable isotopes along the s-process path.
-The seminal work of Goriely Goriely99 introduced a multi-event s-process model to quantify how uncertainties in nuclear reaction rates propagate into s-process abundance predictions and, consequently, into r-process residuals.
-Building on this framework and incorporating the updated s-process calculations of Bisterzo14, recent galactic chemical evolution models have provided refined estimates of s-process contributions across the nuclear chart.
+[…]
+The seminal work of Goriely Goriely (1999) introduced a multi-event s-process model to quantify how uncertainties in nuclear reaction rates propagate into s-process abundance predictions and, consequently, into r-process residuals.
+Building on this framework and incorporating the updated s-process calculations of Bisterzo et al.
+(2014), recent galactic chemical evolution models have provided refined estimates of s-process contributions across the nuclear chart.
 Nevertheless, significant uncertainties remain for several key isotopes, where large MACS uncertainties translate directly into poorly constrained r-process abundances.
@@
 Many of these isotopes are located at critical positions along the s-process path—such as branching points or neutron magic numbers—where accurate cross section measurements not only reduce r-process uncertainties but also provide important constraints on s-process physics, including neutron densities, temperatures, and mixing processes in AGB stars.
-From a technical perspective, the neutron-capture measurements at n_TOF employ C6D6 liquid scintillation detectors in combination with the Pulse Height Weighting Technique (PHWT) Moxon63; Macklin67.
+From a technical perspective, the neutron-capture measurements at n_TOF employ C6D6 liquid scintillation detectors in combination with the Pulse Height Weighting Technique (PHWT) Moxon and Rae (1963); Macklin and Gibbons (1967).
 This method, which requires low-efficiency $\gamma$-ray detectors and detailed Monte Carlo simulations of the experimental setup to derive energy-dependent weighting functions, has been extensively validated at n_TOF.
-Dedicated systematic studies Abbondanno04; Domingo04 have demonstrated that the PHWT, when applied with detailed Monte Carlo (MC)-based weighting functions tailored to each specific sample-detector configuration, can achieve reaction yield determinations with an accuracy of 2% or better.
+Dedicated systematic studies Abbondanno et al.
+(2004); Domingo-Pardo (2004) have demonstrated that the PHWT, when applied with detailed Monte Carlo (MC)-based weighting functions tailored to each specific sample-detector configuration, can achieve reaction yield determinations with an accuracy of 2% or better.
 This validation, performed using samples of various dimensions and compositions, confirmed that the technique properly accounts for the influence of sample geometry and surrounding materials on the detector response, which is a critical requirement for high-precision cross-section measurements.
-These developments now enable systematic cross-section determinations at n_TOF with enhanced accuracy, directly addressing the need for improved nuclear data to constrain r-process residuals and refining s-process models of nucleosynthesis Domingo23b; Domingo25.
+These developments now enable systematic cross-section determinations at n_TOF with enhanced accuracy, directly addressing the need for improved nuclear data to constrain r-process residuals and refining s-process models of nucleosynthesis Domingo-Pardo et al.
+(2023, 2025).
 2 Methodology
@@
 To this aim we utilize three primary data sources:
-  • MACS values and uncertainties: The Karlsruhe Astrophysical Database of Nucleosynthesis in Stars (KADoNiS) version 1.0 kadonisv1, which provides recommended Maxwellian-averaged cross sections at $kT=30$ keV along with their uncertainties for isotopes relevant to s-process nucleosynthesis.
-  • Solar system abundances: The updated solar abundance compilation by Lodders et al. (2015) Lodders2025.
+  • MACS values and uncertainties: The Karlsruhe Astrophysical Database of Nucleosynthesis in Stars (KADoNiS) version 1.0 Dillmann (2014), which provides recommended Maxwellian-averaged cross sections at $kT=30$ keV along with their uncertainties for isotopes relevant to s-process nucleosynthesis.
+  • Solar system abundances: The updated solar abundance compilation by Lodders et al. (2015) Lodders et al. (2025).
   • S-process contributions: The galactic chemical evolution calculations of Bisterzo et al.
-(2014) Bisterzo14, which provide predictions for the s-process contribution to each isotope’s solar abundance based on detailed AGB stellar models with updated 13C-pocket structures and neutron capture networks.
+(2014) Bisterzo et al.
+(2014), which provide predictions for the s-process contribution to each isotope’s solar abundance based on detailed AGB stellar models with updated 13C-pocket structures and neutron capture networks.
 The r-process residual abundance for a given isotope is obtained by subtracting the s-process contribution from the total solar abundance:
@@
 Figure 1: Isotopic abundances in the solar system (normalized to Si $10^{6}$) showing the decomposition into s-process and r-process components.
-Upper panel: Total solar system abundances from Lodders2025 (open circles), s-process contributions from Bisterzo14 (blue filled circles), and derived r-process residuals $N_{\rm r}=N_{\odot}-N_{\rm s}$ (red filled circles).
+Upper panel: Total solar system abundances from Lodders et al.
+(2025) (open circles), s-process contributions from Bisterzo et al.
+(2014) (blue filled circles), and derived r-process residuals $N_{\rm r}=N_{\odot}-N_{\rm s}$ (red filled circles).
 Lower panel: Relative s-process contributions (percentage of total solar abundance).
@@
 Although this approximation neglects second-order effects and correlations between isotopes in network calculations, it provides a practical first-order estimate for identifying the most problematic cases.
-The approximation is well justified by the s-process equilibrium conditions in AGB stars: in the mass range A=90A=90–200, the product $N_{s}\langle\sigma\rangle$ remains approximately constant Kaeppeler11, which directly supports the assumed linear scaling between MACS uncertainties and s-process abundance uncertainties.
-Given that uncertainties in solar system abundances from Lodders2025 are typically small ($\sim$1–3%) compared to the dominant MACS uncertainties (ranging from $\sim$10% to >>50% for many isotopes), we approximate the uncertainty in the r-process residual as:
+The approximation is well justified by the s-process equilibrium conditions in AGB stars: in the mass range A=90A=90–200, the product $N_{s}\langle\sigma\rangle$ remains approximately constant Käppeler et al.
+(2011), which directly supports the assumed linear scaling between MACS uncertainties and s-process abundance uncertainties.
+We note that our analysis focuses exclusively on the main s-process component produced in AGB stars.
+The weak s-process component, follows different nucleosynthesis conditions Pignatari et al.
+(2010) and is not addressed in the current work.
+Nevertheless, the methodology presented here could be extended to that mass region in future studies, where uncertainties in r-process residuals are expected to be similar or even more significant.
+Given that uncertainties in solar system abundances from Lodders et al.
+(2025) are typically small ($\sim$1–3%) compared to the dominant MACS uncertainties (ranging from $\sim$10% to >>50% for many isotopes), we approximate the uncertainty in the r-process residual as:
 $\Delta N_{\rm r}\approx\Delta N_{\rm s}=N_{\rm s}\cdot\frac{\Delta\sigma}{\langle\sigma\rangle}$ (3)
-This simplified approach allows us to directly identify isotopes where improved MACS measurements would have the greatest impact on constraining r-process yields through reduction of uncertainties in the derived r-process residuals.
-The resulting r-process abundances are shown in Fig.1.
-As it can be inferred from Eq.3, r-process residual uncertainties become larger the higher the s-process contribution (Ns) and the higher the relative uncertainty on the MACS.
+This simplified approach provides a practical lower limit on the uncertainty in r-process residuals, as it assumes MACS uncertainties dominate over other potential error sources such as uncertainties in s-process stellar models, beta-decay rates at branching points, and temperature-dependent effects.
+While this assumption is well justified for many isotopes where MACS uncertainties exceed 10–15% and solar abundance uncertainties remain below 3%, there are cases, particularly at certain branching points, where additional uncertainty contributions become significant (see discussion of 99Ru in Section 3).
+Nevertheless, this first-order approximation successfully identifies isotopes where improved MACS measurements would have substantial impact on constraining r-process yields.
+A more comprehensive uncertainty analysis incorporating all error sources through full s-process network calculations would be valuable but is beyond the scope of this prioritization study.
 For this study we identify isotopes as high-priority targets for improved $(n,\gamma)$ measurements based on two criteria:
@@
 200Hg $4.3(9)\times 10^{-2}$ 22 10
-Table 1: Isotope data with Nr values, uncertainties, and relative uncertainties for Nr and MACS.
+Table 1: Isotope data with Nr values, uncertainties, and relative uncertainties for $N_{r}$ and MACS.
 3 Case Studies of Key Isotopes
 97Mo
-Molybdenum isotopes serve as critical tracers of s-process nucleosynthesis in presolar silicon carbide grains from AGB stars Stephan2019.
-While 96Mo is especially valuable as an s-process-only isotope, shielded from r-process production by 96Zr, 97Mo is a mixed s- and r-process isotope contributing about 47% to the solar inventory via s-process.
-The high-precision measurements of these isotopes in mainstream SiC grains Liu24 reveal that their compositions follow clear mixing lines between a pure s-process progenitor and solar material, with 96Mo and 97Mo abundances closely following the inverse of their neutron-capture cross sections.
-This behavior shows remarkably constant s-process ratios across different SiC grains, indicating that the branching at 95Zr remains minimal even when the 22Ne($\alpha$,n)25Mg reaction is marginally activated during thermal pulses at peak neutron densities of $\sim 10^{10}$ cm-3 Stephan2019; Lugaro03; Liu2014.
+Molybdenum isotopes serve as critical tracers of s-process nucleosynthesis in presolar silicon carbide grains from AGB stars Stephan et al.
+(2019).
+While 96Mo is especially valuable as an s-process-only isotope (see Fig.2), shielded from r-process production by 96Zr, 97Mo is a mixed s- and r-process isotope contributing about 47% to the solar inventory via s-process.
+The high-precision measurements of these isotopes in mainstream SiC grains Liu (2025) reveal that their compositions follow clear mixing lines between a pure s-process progenitor and solar material, with 96Mo and 97Mo abundances closely following the inverse of their neutron-capture cross sections.
+This behavior shows remarkably constant s-process ratios across different SiC grains, indicating that the branching at 95Zr remains minimal even when the 22Ne($\alpha$,n)25Mg reaction is marginally activated during thermal pulses at peak neutron densities of $\sim 10^{10}$ cm-3 Stephan et al.
+(2019); Lugaro et al.
+(2003); Liu et al.
+(2014).
 This weak branching is sufficient to produce the observed 96Zr abundances without significantly bypassing 96Mo in favor of 97Mo and 98Mo.
 The observed variations in other Mo isotopes like 94Mo, 95Mo, and 100Mo provide unique constraints on neutron densities, temperatures, and timescales during s-process nucleosynthesis in the parent AGB stars.
-Recent measurements of the 96Mo neutron capture cross section have been performed at the n_TOF facility (CERN) and complementary transmission measurements at GELINA (JRC-Geel) Mucciola22; Mucciola23.
+Figure 2: Nuclear chart segment showing the 99Tc branching point that determines 99Ru production in AGB stars.
+Recent measurements of the 96Mo neutron capture cross section have been performed at the n_TOF facility (CERN) and complementary transmission measurements at GELINA (JRC-Geel) Mucciola et al.
+(2022, 2023).
 The measurements utilized isotopically enriched 96Mo samples with 95.90% enrichment.
-Capture cross sections were measured at both the EAR2 (18 m from spallation target) and EAR1 (184 m) experimental stations of n_TOF during the 2021–2022 campaigns, using arrays of C6D6 liquid scintillation detectors and newly developed sTED detectors Alcayne24; Balibrea25.
+Capture cross sections were measured at both the EAR2 (18 m from spallation target) and EAR1 (184 m) experimental stations of n_TOF during the 2021–2022 campaigns, using arrays of C6D6 liquid scintillation detectors and newly developed sTED detectors Alcayne et al.
+(2024); Balibrea-Correa et al.
+(2025).
 Complementary transmission measurements were performed at the 10 m station of GELINA using a 6Li-glass neutron detector.
@@
 These measurements are expected to significantly reduce the current 14% MACS uncertainty.
-For 97Mo, no experimental neutron-capture cross section measurement exists to date and an experiment proposal has been approved for a high-resolution high-accuracy measurement at CERN n_TOF Busso2024.
-Given the similar nuclear structure and expected cross section magnitude compared to 96Mo, a measurement achieving comparable precision (reducing the current 15% MACS uncertainty kadonisv1 to $\sim$4–5%) is feasible.
+For 97Mo, no experimental neutron-capture cross section measurement exists to date and an experiment proposal has been approved for a high-resolution high-accuracy measurement at CERN n_TOF Busso et al.
+(2024).
+Given the similar nuclear structure and expected cross section magnitude compared to 96Mo, a measurement achieving comparable precision (reducing the current 15% MACS uncertainty Dillmann (2014) to $\sim$4–5%) is feasible.
 This would substantially improve the determination of the r-process residual, currently uncertain by $\sim$27%.
@@
 Given that theoretical predictions more typically carry uncertainties as high as a factor of 2 even for nuclei close to stability, the quoted 15% uncertainty is almost certainly underestimated.
-Moreover, this is one of the very few stable isotopes where no direct neutron-capture measurement exists in the stellar energy range.
-The apparent moderate uncertainty on the r-process residual (32%) may be a consequence of the optimistic uncertainty assigned to the unknown neutron-capture cross section.
+Moreover, the total uncertainty in the s-process production of 99Ru is not determined solely by the 99Tc($n,\gamma$) cross section but also by the stellar $\beta$-decay rate of 99Tc.
+Uncertainties in stellar decay rates at this branching point can reach factors of 5-10 due to thermal population of excited states and temperature dependency Schatz (1983); Cosner and Truran (1981); Takahashi and Yokoi (1987).
+For 99Tc, the competition between neutron capture and $\beta$-decay determines the branching ratio and thus the final 99Ru abundance.
+Both nuclear physics inputs contribute to the total uncertainty budget.
+Therefore, while a precision measurement of the 99Ru($n,\gamma$) cross section would reduce one component of the uncertainty, a complete understanding of the 99Ru production in AGB stars requires both, improved accuracy on the 99Ru($n,\gamma$) cross section, and better constraints on the 99Tc stellar $\beta$-decay rate.
+The current $\sim$32% uncertainty on the r-process residual of 99Ru likely underestimates the true uncertainty when the branching point effect is fully considered.
+Nevertheless, measuring the 99Ru($n,\gamma$) cross section remains a high-priority experiment as it addresses a fundamental gap in our nuclear data-set (this is one of the few stable isotopes with no direct measurement) and provides essential input for the comprehensive s-process network calculations.
 With a natural abundance of 12.76% and a large expected cross section of 631 mb at $kT=30$ keV, producing an enriched sample of 99Ru for a time-of-flight measurement should be straightforward.
-Reducing the uncertainty to the 4-5% level should lead to an r-process residual accurate to 6%.
-From the viewpoint of the s-process, this measurement would be extremely valuable for a fully consistent interpretation of the first detection of extinct 99Tc in presolar silicon carbide grains, which was achieved by measuring an anomaly in its decay product 99Ru using resonance ionization mass spectrometry Savina2004.
+Considering only the cross-section uncertainty, by reducing it to the 4-5% level should lead to an r-process residual accurate to 6%.
+From the viewpoint of the s-process, this measurement would be extremely valuable for a fully consistent interpretation of the first detection of extinct 99Tc in presolar silicon carbide grains, which was achieved by measuring an anomaly in its decay product 99Ru using resonance ionization mass spectrometry Savina et al.
+(2004).
 The isotopic composition of ruthenium in individual SiC stardust grains bears the signature of s-process nucleosynthesis in asymptotic giant branch stars, with an enrichment in 99Ru explained by the in situ decay of 99Tc that was present when the grains formed.
-This finding directly connects to Merrill’s 1952 observation of technetium spectral lines in S-type stars Merrill52, which provided the first evidence that elements are synthesized within stars.
+This finding directly connects to Merrill’s 1952 observation of technetium spectral lines in S-type stars Merrill (1952), which provided the first evidence that elements are synthesized within stars.
 By detecting the radiogenic 99Ru signature left by now-extinct 99Tc, Savina et al.
-Savina2004 confirmed that the majority of presolar SiC grains originated from low-mass AGB stars and established that the amount of 99Tc produced in these stars is insufficient to leave a detectable 99Ru anomaly in bulk early solar system materials.
+Savina et al.
+(2004) confirmed that the majority of presolar SiC grains originated from low-mass AGB stars and established that the amount of 99Tc produced in these stars is insufficient to leave a detectable 99Ru anomaly in bulk early solar system materials.
 This work allowed stellar nucleosynthesis models to be quantitatively tested with unprecedented accuracy by bridging spectroscopic observations of active stellar nucleosynthesis with laboratory measurements of preserved stellar ejecta.
 122Sn
-The pioneering work of Macklin and co-workers in 1962 Macklin62 demonstrated that tin’s ten stable isotopes provide exceptional opportunities to disentangle the contributions of different neutron-capture processes.
-According to the original hypothesis of heavy-element nucleosynthesis BBFH, certain tin isotopes can only be produced through specific pathways: 116Sn cannot form via r-process, while 122Sn and 124Sn were thought to be exclusive products of the rapid neutron capture process.
-By measuring neutron capture cross-sections of the tin isotopes at stellar temperatures and examining the inverse relationship between these cross-sections and isotopic abundances it was possible to quantify for the first time the relative contributions of s-process versus r-process nucleosynthesis in the solar system’s material Macklin62.
+The pioneering work of Macklin and co-workers in 1962 Macklin et al.
+(1962) demonstrated that tin’s ten stable isotopes provide exceptional opportunities to disentangle the contributions of different neutron-capture processes.
+According to the original hypothesis of heavy-element nucleosynthesis Burbidge et al.
+(1957), certain tin isotopes can only be produced through specific pathways: 116Sn cannot form via r-process, while 122Sn and 124Sn were thought to be exclusive products of the rapid neutron capture process (see Fig.3).
+By measuring neutron capture cross-sections of the tin isotopes at stellar temperatures and examining the inverse relationship between these cross-sections and isotopic abundances it was possible to quantify for the first time the relative contributions of s-process versus r-process nucleosynthesis in the solar system’s material Macklin et al.
+(1962).
 This approach confirmed theoretical predictions about how these processes operate in different stellar environments: the s-process occurring in giant stars and the r-process in explosive environments.
 The approximate constancy of the product of neutron capture cross-section and abundance for s-process isotopes provided direct experimental validation of nucleosynthesis models, establishing tin as an ideal “laboratory” element for testing our understanding of how heavy elements form in stars.
-Today, thanks to detailed AGB models we know that 122Sn is not an exclusive product of the r-process Bisterzo14.
+Figure 3: Tin isotopes spanning the transition from s-process-only 116Sn (not shown) to mixed r/s-process nuclei 122Sn and 124Sn.
+Today, thanks to detailed AGB models we know that 122Sn is not an exclusive product of the r-process Bisterzo et al.
+(2014).
 However, despite the critical role of 122Sn in disentangling s- and r-process contributions in this mass region, experimental constraints on its neutron capture cross section remain limited.
-The current nuclear data situation relies primarily on a single time-of-flight measurement by Timokhov1989 and two activation measurements Stadler1998; Murty1973, one of them unpublished.
+The current nuclear data situation relies primarily on a single time-of-flight measurement by Timokhov et al.
+(1989) and two activation measurements Stadler (1998); Murty et al.
+(1973), one of them unpublished.
 The recommended MACS carries an uncertainty of 35%, which propagates to an r-process residual uncertainty exceeding 50% when performing classical s/r-process decomposition.
-This large uncertainty significantly hampers our ability to use 122Sn as a diagnostic tool for understanding the branching-point physics and for constraining galactic chemical evolution models Bisterzo14.
+This large uncertainty significantly hampers our ability to use 122Sn as a diagnostic tool for understanding the branching-point physics and for constraining galactic chemical evolution models Bisterzo et al.
+(2014).
 Given its relatively high natural abundance of 4.63%, 122Sn represents an excellent candidate for high-precision measurements combining time-of-flight and activation techniques.
-New measurements at the CERN n_TOF facility, utilizing both the EAR1 and EAR2 beam lines as well as the NEAR station for activation studies Patronis25, would provide the complementary data needed to substantially reduce these uncertainties and better constrain the s-process contribution to this key branching-point isotope.
+New measurements at the CERN n_TOF facility, utilizing both the EAR1 and EAR2 beam lines as well as the NEAR station for activation studies N.
+Patronis et al.
+n_TOF Collaboration (2025), would provide the complementary data needed to substantially reduce these uncertainties and better constrain the s-process contribution to this key branching-point isotope.
 139La
-139La is the most abundant isotope of lanthanum (>99.9>99.9%) and is primarily produced in the s-process, with an estimated contribution of 75% Bisterzo14.
-It is of particular interest since it has a magic neutron number ($N=82$) and as a consequence a small neutron capture cross section compared to neighbouring isotopes.
+139La is the most abundant isotope of lanthanum (>99.9>99.9%) and is primarily produced in the s-process, with an estimated contribution of 75% Bisterzo et al.
+(2014).
+It is of particular interest since it has a magic neutron number ($N=82$) and as a consequence a small neutron capture cross section compared to neighbouring isotopes (see Fig.4).
 Hence, the 139La($n,\gamma$) reaction represents a bottleneck in the s-process flow where matter accumulates.
 Based on the current uncertainties of the Maxwellian averaged cross section of around 10%, the resultant uncertainty of the solar r-process abundance is around 39%.
-There are several measurements of the 139La($n,\gamma$) cross section kadonisv1.
-While available experimental MACS values at $kT=30$ keV range from 22.2 to 50 mb Dillmann08, the three most recent results quote consistent values of $31.6\pm 0.8$ mb Obrien03, $31.2$ mb (renormalised from a measurement at $kT=5$ keV) Winckler_2006 and $32.4\pm 2.1$ mb Terlizzi07.
-The recommended value in the KADoNiS database and hence the value most frequently used in stellar model calculations is the result from measurement at n_TOF by Terlizzi et al Terlizzi07.
+Figure 4: Lanthanum isotopes near the N = 82 magic number, where 139La acts as an s-process bottleneck.
+There are several measurements of the 139La($n,\gamma$) cross section Dillmann (2014).
+While available experimental MACS values at $kT=30$ keV range from 22.2 to 50 mb Dillmann et al.
+(2008), the three most recent results quote consistent values of $31.6\pm 0.8$ mb O’Brien et al.
+(2003), $31.2$ mb (renormalised from a measurement at $kT=5$ keV) Winckler et al.
+(2006) and $32.4\pm 2.1$ mb Terlizzi et al.
+(2007).
+The recommended value in the KADoNiS database and hence the value most frequently used in stellar model calculations is the result from measurement at n_TOF by Terlizzi et al Terlizzi et al.
+(2007).
 This measurement covered neutron energy ranges from 0.6 eV to 9 keV.
@@
 The activation measurement by O’Brien et al.
-Obrien03 is the only direct measurement at 30 keV neutron energy.
+O’Brien et al.
+(2003) is the only direct measurement at 30 keV neutron energy.
 The quoted uncertainty is small and does not include the uncertainty of the 197Au($n,\gamma$) the spectrum averaged cross section used as a reference reaction.
-Incidentally, the 197Au($n,\gamma$) in the keV neutron energy region has recently been updated Lederer11; Massimi14; Reifarth2018, which would lead to a small reduction of the value obtained by Obrien03.
-Concerning the n_TOF measurement Terlizzi07, the energy range analysed was limited by the increase of in-beam $\gamma$-ray background in the keV neutron energy range, combined with low statistics due to the small neutron capture cross section.
+Incidentally, the 197Au($n,\gamma$) in the keV neutron energy region has recently been updated Lederer et al.
+(2011); Massimi et al.
+(2014); Reifarth et al.
+(2018), which would lead to a small reduction of the value obtained by O’Brien et al.
+(2003).
+Concerning the n_TOF measurement Terlizzi et al.
+(2007), the energy range analysed was limited by the increase of in-beam $\gamma$-ray background in the keV neutron energy range, combined with low statistics due to the small neutron capture cross section.
 Since then, the spallation target at n_TOF was replaced, and now borated water is used as neutron moderator.
-This results in a significant reduction of the in-beam $\gamma$-ray background (see Fig.13 in Ref.Massimi10, and Refs.
-Lerendegui16a; Lerendegui16b), and neutron capture cross sections have since been routinely measured up to tens or hundreds of keV, even for comparably small cross sections, e.g.
-Lederer14; Gawlik21; Amaducci24.
+This results in a significant reduction of the in-beam $\gamma$-ray background (see Fig.13 in Ref.Massimi et al.
+(2010), and Refs.
+Lerendegui-Marco et al.
+(2016a, b)), and neutron capture cross sections have since been routinely measured up to tens or hundreds of keV, even for comparably small cross sections, e.g.
+Lederer et al.
+(2014); Gawlik-Ramiega et al.
+(2021); Amaducci et al.
+(2024).
 A new measurement, aiming for high counting statistics at n_TOF would allow to access the entire neutron energy range relevant for AGB stars (the main producers of the s-process component of 139La), and reduce uncertainties to 5-6%, typically achieved at n_TOF.
 This can be combined with an activation measurement at n_TOF NEAR.
-In parallel, a re-normalisation of the O’Brien Obrien03 values based on the updated 197Au($n,\gamma$) cross section will help to provide a second independent value for the MACS at $kT=30$ keV.
+In parallel, a re-normalisation of the O’Brien O’Brien et al.
+(2003) values based on the updated 197Au($n,\gamma$) cross section will help to provide a second independent value for the MACS at $kT=30$ keV.
 168Er
-Erbium isotopes, including 168Er, occupy a critical region in the landscape of s-process nucleosynthesis.
+Erbium isotopes, including 168Er, occupy a critical region in the landscape of s-process nucleosynthesis (see Fig.5).
 Located beyond the N = 82 neutron magic number in the mass range A $\sim$ 166–170, the erbium isotopes lie in the transition zone between the second s-process abundance peak (A $\sim$ 140) and the third peak (A $\sim$ 208).
+Figure 5: Erbium isotopes in the deformed mid-shell region between the second and third s-process abundance peaks.
 The s-process dominates the production of elements in the erbium-ytterbium (Er–Yb) region, with contributions of 70–90% of the solar system abundances.
@@
 […]
-The most recent direct measurement Marganiec09 employed the activation technique at the Karlsruhe Van de Graaff accelerator, yielding a MACS at $kT=30$ keV of $221\pm 12$ mb (5.4% uncertainty).
-This value confirmed the previously recommended MACS of $224\pm 10$ mb Bao00 and showed good agreement with earlier time-of-flight measurements Macklin83; Beer82.
-However, when combined with new data for the neighboring isotopes 180W and 186W, the measurements revealed a significantly flatter cross-section trend with mass number than predicted by Hauser-Feshbach calculations Rauscher00; Goriely, highlighting remaining challenges in theoretical modeling and indicating that statistical model predictions commonly used for r-process network calculations may overestimate cross sections for the neutron-rich tungsten isotopes.
-From the viewpoint of enhanced accuracy measurements, improving the presently recommended relative uncertainty of $\sim$12%kadonisv1 down to 4% could still help to reduce the 57% uncertainty on the r-process isotopic abundance down to $\sim$20%.
-Such a measurement at CERN n_TOF could benefit of the availability of both TOF- and activation techniques and new instrumentation at EAR1, EAR2 Alcayne24; Balibrea25 and NEAR Patronis25 stations.
+Figure 6: Tungsten isotopes near the 185W branching point, critical for Re/Os cosmochronometry.
+The most recent direct measurement Marganiec et al.
+(2009) employed the activation technique at the Karlsruhe Van de Graaff accelerator, yielding a MACS at $kT=30$ keV of $221\pm 12$ mb (5.4% uncertainty).
+This value confirmed the previously recommended MACS of $224\pm 10$ mb Bao et al.
+(2000) and showed good agreement with earlier time-of-flight measurements Macklin et al.
+(1983); Beer et al.
+(1982).
+However, when combined with new data for the neighboring isotopes 180W and 186W, the measurements revealed a significantly flatter cross-section trend with mass number than predicted by Hauser-Feshbach calculations Rauscher and Thielemann (2000); Goriely (2007), highlighting remaining challenges in theoretical modeling and indicating that statistical model predictions commonly used for r-process network calculations may overestimate cross sections for the neutron-rich tungsten isotopes.
+From the viewpoint of enhanced accuracy measurements, improving the presently recommended relative uncertainty of $\sim$12%Dillmann (2014) down to 4% could still help to reduce the 57% uncertainty on the r-process isotopic abundance down to $\sim$20%.
+Such a measurement at CERN n_TOF could benefit of the availability of both TOF- and activation techniques and new instrumentation at EAR1, EAR2 Alcayne et al.
+(2024); Balibrea-Correa et al.
+(2025) and NEAR N.
+Patronis et al.
+n_TOF Collaboration (2025) stations.
 A measurement which, considering the relatively large cross section and natural abundance of 184W (30%) should be rather straight-forward.
 200Hg
-The r-process residual derived from 200Hg may serve as an important anchor point for constraining the r-process abundance distribution in the $A\sim 200$ mass region.
+The r-process residual derived from 200Hg may serve as an important anchor point for constraining the r-process abundance distribution in the $A\sim 200$ mass region (see Fig.7).
 Together with the other mercury isotopes (199,201,202Hg), the 200Hg data help to define the relatively flat r-process abundance plateau expected in this mass range, lying between the $N=126$ r-process peak and the lead region.
 This information is helpful for validating r-process nucleosynthesis models and for understanding the conditions under which rapid neutron capture occurred in the early Galaxy.
+[…]
 With a measured MACS of $115\pm 12$ mb at $kT=30$ keV, 200Hg represents a key (s+r) process isotope that receives rather balanced contributions from both slow and rapid neutron-capture processes.
 Thus, the accurate cross section determination enables quantitative decomposition of the 200Hg solar abundance into its s- and r-process components.
-Using the measured cross section and the $\sigma N$ systematics of the s-process, Beer and Macklin determined that of the total solar abundance of $(0.045\pm 0.007)/10^{6}$ Si, the s-process contributes 42(5)% while the r-process accounts for 55(16)%Beer85.
+Using the measured cross section and the $\sigma N$ systematics of the s-process, Beer and Macklin determined that of the total solar abundance of $(0.045\pm 0.007)/10^{6}$ Si, the s-process contributes 42(5)% while the r-process accounts for 55(16)%Beer and Macklin (1985).
 A new measurements could significantly decrease the uncertainty, down to $\sim$5%, thus lowering also the relative r-process uncertainty down to the level of $\sim$10%.
@@
 We have presented a systematic analysis of neutron capture cross section uncertainties and their impact on r-process residual determinations for isotopes in the mass range $A\sim 96$–200.
-By combining MACS data from KADoNiS 1.0 kadonisv1, solar system abundances from Lodders2025, and s-process predictions from Bisterzo14, we have identified a set of high-priority targets where improved measurements would significantly reduce uncertainties in derived r-process yields.
+By combining MACS data from KADoNiS 1.0 Dillmann (2014), solar system abundances from Lodders et al.
+(2025), and s-process predictions from Bisterzo et al.
+(2014), we have identified a set of high-priority targets where improved measurements would significantly reduce uncertainties in derived r-process yields.
 Our analysis reveals that seven isotopes meet the dual criteria of MACS relative uncertainty $\geq 10\%$ and r-process residual relative uncertainty $\geq 20\%$: 97Mo, 99Ru, 122Sn, 139La, 168Er, 184W and 200Hg.
@@
 The enhanced capabilities of CERN n_TOF following Long Shutdown 3 present an excellent opportunity to address these measurement needs.
-The combination of the high-flux EAR1 and EAR2 Pavon25 beam lines with the newly commissioned NEAR station Patronis25 provides optimal conditions for both time-of-flight and activation measurements across the relevant energy range.
-The improved neutron flux, enhanced detection systems Alcayne24; Balibrea25, and extended measurement capabilities will enable precision determinations of MACS values with target uncertainties below 5% for most of the identified isotopes.
+The combination of the high-flux EAR1 and EAR2 Pavon-Rodriguez et al.
+(2025) beam lines with the newly commissioned NEAR station N.
+Patronis et al.
+n_TOF Collaboration (2025) provides optimal conditions for both time-of-flight and activation measurements across the relevant energy range.
+The improved neutron flux, enhanced detection systems Alcayne et al.
+(2024); Balibrea-Correa et al.
+(2025), and extended measurement capabilities will enable precision determinations of MACS values with target uncertainties below 5% for most of the identified isotopes.
 Beyond the immediate goal of constraining r-process residuals, these measurements will provide critical input for s-process nucleosynthesis models themselves.
@@
 Acknowledgments
-We dedicate this work to the memory of Roberto Gallino, whose pioneering theoretical work on s-process nucleosynthesis in AGB stars provided the essential framework for interpreting neutron-capture measurements such as those performed at FZK Karlsruhe and CERN n_TOF Gallino23.
+We dedicate this work to the memory of Roberto Gallino, whose pioneering theoretical work on s-process nucleosynthesis in AGB stars provided the essential framework for interpreting neutron-capture measurements such as those performed at FZK Karlsruhe and CERN n_TOF Gallino (2023).
 His detailed stellar models, incorporating the physics of 13C-pockets and nuclear-physics networks, have been instrumental in calculating the s-process contributions needed to understand the composition of our solar system and the chemical evolution of our galaxy.
@@
 Dillmann for facilitating quick access to the data of their publications and evaluations.
+Fig.
+2 to Fig.
+7 have been adapted from the KADoNiS web database.
 The authors acknowledge support from all the funding agencies of participating institutions.